OpenStudy (anonymous):

Find f^(-1) when f(x)=tanhx

OpenStudy (anonymous):

You know that\[\sinh(x)=\frac{e^{x}-e^{-x}}{2}, \cosh(x)=\frac{e^{x}+e^{-x}}{2}\] So\[f(x)=\frac{e^{2x}-1}{e^{2x}+1}\] here I denote \[s=\frac{e^{2x}-1}{e^{2x}+1}\] How do you isolate x though? I'm stuck. :(